OpenStudy (anonymous):

show that (cos^4theta-sin^4theta)/cos^2theta=1-tan^2theta

OpenStudy (anonymous):

(cos[theta]+tan[theta])/(sin[theta]) cos[theta]/sin[theta] + tan[theta]/sin[theta] cot[theta] + 1/cos[theta] cot[theta] + sec[theta]

OpenStudy (anonymous):

that doesn't prove it

OpenStudy (campbell_st):

1st factorise the numerator (\[\frac{(\cos^2\theta + \sin^2\theta)(\cos^2\theta - \sin^2\theta)}{\cos^2\theta}\] use \[\cos^2\theta + \sin^2 \theta = 1\] then the problem is \[\frac{1\times(\cos^2\theta - \sin^2\theta)}{\cos^2\theta}\] then splitting the terms in the numerator \[\frac{\cos^2\theta}{\cos^2 \theta} - \frac{\sin^2\theta}{\cos^2\theta}\] using \[\tan \theta =\frac{\sin \theta}{\cos \theta}\] gives \[1 - \tan^2 \theta\]